# M3DLayout: A Multi-Source Dataset of 3D Indoor Layouts and Structured Descriptions for 3D Generation

### v1

@@ 3.3 Dataset Statistics and Analysis
Data Composition and Source Characteristics. As detailed in Table 1, M3DLayout integrates 15,080 layouts with 258,413 object instances, averaging 17.1 objects per scene. The three data sources exhibit complementary characteristics: 3D-FRONT provides professionally designed layouts with clean structural regularity but limited small objects (0.2%); Matterport3D offers realistic scanned environments with moderate object density (12.6 objects/scene) and a balanced object distribution (39.4% small objects); while Inf3DLayout significantly enriches the dataset with high scene complexity (25.9 objects/scene) and abundant small objects (71.1%).
Comparative Advantages. As shown in Table 2, M3DLayout surpasses existing datasets in scale (15,080 scenes, 258k+ objects), layout complexity, and object size variation. Unlike datasets limited to either large furniture (L) or simple layouts, M3DLayout provides comprehensive coverage of both large and small objects (L-S) with structured textual descriptions, addressing a critical gap in current data resources for detailed scene generation.
4 Benchmark
To evaluate the capabilities and limitations of the M3DLayout dataset, we establish a comprehensive benchmark using a diffusion model to assess the dataset’s potential in supporting text-conditioned 3D indoor layout generation.
4.1 Problem Formulation
We formulate 3D indoor layout generation as a conditional denoising diffusion process. Given a scene-level natural language description $c^{\text{text}}$, our goal is to generate a structured 3D layout $x$ consisting of $N$ objects. Each object $o_{i}$ in the layout is parameterized as a 3D oriented bounding box $o_{i}=(c_{i},x_{i},y_{i},z_{i},w_{i},h_{i},d_{i},\theta_{i}),$ where $c_{i}$ denotes the semantic class label, $(x_{i},y_{i},z_{i})$ is the object center in 3D space, $(w_{i},h_{i},d_{i})$ represent the width, height, and length, and $\theta_{i}$ is the yaw angle. The full layout is then denoted as $x_{0}=\{o_{i}\}_{i=1}^{N}.$
Following the Denoising Diffusion Probabilistic Model (DDPM) framework, we define a forward diffusion process that gradually adds Gaussian noise to the layout representation $q(x_{t}\mid x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I}),$ where $\beta_{t}$ is a predefined noise schedule. The reverse process is modeled by a neural network $\epsilon_{\theta}$ that predicts the noise $\epsilon$ from the noisy layout $x_{t}$ and the conditioning information $c^{\text{text}}$:
The model is trained to minimize the noise prediction objective $\mathcal{L}_{\text{DM}}=\mathbb{E}_{x_{0},c^{\text{text}},t,\epsilon}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t,c^{\text{text}})\|_{2}^{2}\right].$
4.2 Model Architecture
Our model architecture follows a denoising diffusion framework similar to DiffuScene, employing a U-Net style backbone with 1D convolutional layers and attention mechanisms, with conditional inputs injected via cross-attention.
For the training objective, we employ the standard noise prediction loss combined with additional regularization terms. The primary diffusion loss minimizes the L2 distance between predicted and actual noise components. We also incorporate an IoU loss term to penalize object intersections and encourage physically plausible arrangements.
4.3 Experimental Settings
We conduct experiments to evaluate the effectiveness of our conditional diffusion model on the proposed dataset. The goal is to assess both the plausibility and controllability of the generated 3D layouts under different scene conditions. The detailed dataset train/val splits and implementation details are provided in the Appendix A.1. We describe our evaluation settings as follows:
Baselines. We compare our method with two state-of-the-art scene generation approaches: DiffuScene Tang et al. (2024) and InstructScene Lin & Mu (2024), both of which are text-driven methods based on diffusion models. More details are provided in the Appendix A.2.
Metrics. Following prior works Tang et al. (2024); Lin & Mu (2024), we adopt Fréchet Inception Distance (FID) Heusel et al. (2017) and Kernel Inception Distance (KID) Bińkowski et al. (2018), and the CLIP-Score Hessel et al. (2021) to measure the controllability and fidelity, respectively. We employ a self-designed object retrieval method to realize instance filling from layout to scene and rendering. More details about rendering and retrieval are provided in the Appendix A.3 and Appendix A.6.
Table 3: Quantitative comparison of layout generation methods and ablation studys of our model trained on different datasets. Lower FID/KID ($\times$0.001) and higher Clip Score indicate better synthesis quality. FID and KID are computed with respect to the real layouts from 3D-FRONT, Matterport, and Inf3DLayout. We train InstructScene following the public implementation.
@@ 4.3 Experimental Settings
DiffuSceneTang et al. (2024) 29.47 98.03 102.12 10.32 47.92 75.49 0.1982
InstructSceneLin & Mu (2024) 68.58 100.54 159.27 54.70 49.23 156.62 0.1944
Ours (M3DLayout) 57.64 87.89 70.85 36.80 34.62 50.94 0.2001
Ours (3D-FRONT) 27.33 83.88 110.98 10.59 21.80 83.45 0.2083
@@
4.4 Quantitative and Qualitative Comparison
Quantitative Comparison. The quantitative comparison results are shown in the upper part of Table 3, where DiffuScene and InstructScene are trained on 3D-FRONT and Ours on M3DLayout. In Table 3, the horizontal axis represents ground-truth renderings (real images) from three different datasets. As described in Appendix A.1, the same set of 1,500 prompts is used as conditioning input for all comparative methods to generate layouts, which are then rendered as synthetic images and used together with real images to compute FID and KID. Typically, real images are taken from the training set; however, since our method and the baselines are trained on different datasets with varying data distributions, direct comparison would be unfair. To address this, we select different datasets as the source of real images, allowing for both a fair comparison and an evaluation of fidelity and controllability.
As demonstrated, for FID and KID, our method drastically outperforms the state-of-the-art by 10%–32% on the reference Matterport and Inf3DLayout dataset, demonstrating superior generalization compared with DiffuScene and InstructScene. On 3D-FRONT, however, it falls behind DiffuScene on these metrics. This is primarily because the number of objects per scene in 3D-FRONT typically ranges from 5 to 12, whereas our method generates scenes with more than 12 objects in most cases, leading to a mismatch in the distribution of scene complexity. As a result, the visual statistics of our generated layouts deviate more from the ground truth (3D-FRONT), which negatively impacts FID and KID. Conversely, the richer object counts and variety in the scenes generated by our model provide another advantage of our method relative to the multi-source datasets M3DLayout, which is also confirmed by the visualizations in Figure 5. Moreover, our method surpasses the baselines in CLIP-Score, demonstrating enhanced controllability and a stronger alignment of the generated scenes with the given prompts.
Qualitative Comparison. The qualitative comparison results for different methods in the bedroom, dining room, and living room are shown in Figure 5. Across all settings, our method trained on the M3DLayout dataset demonstrates improved semantic controllability and visual fidelity compared to Diffuscene and Instructscene. Specifically, DiffuScene can generate scenes that appear visually neat at first glance but struggles to produce small objects and exhibits limited prompt controllability, as illustrated by the living room case. InstructScene, on the other hand, fails to accurately model spatial relationships among instances, leading to disorderly object placement, as shown in the dining room case. By contrast, our method generates precise and diverse objects in accordance with the prompt semantics—such as the small items on the shelf and six chairs in the dining room case—while effectively capturing 3D spatial arrangements, exemplified by “The bed is located near the wall corner” in the bedroom case.
4.5 Density Controllability
Figure 4: Density Controllability in Layout Generation with Different Input Prompts. The first row presents input prompts for our layout generation model, showcasing variations in objects density from low to high, with minor changes in the last sentence. The second row illustrates the corresponding output results generated by our model, which adapt based on the prompt density.
4.6 Ablation Study
@@
We perform the ablation experiments to validate the effectiveness of a single training dataset and report the results in the lower part of Table 3. As shown in the table, when the training data and ground truth come from the same dataset, our model trained on a single dataset achieves the best FID and KID compared with the other two methods. However, its performance drops significantly when evaluated on data from different datasets. This indicates that, while models trained on a single dataset can effectively fit the distribution of that dataset, they struggle to generalize to varied data. For example, a model trained on the professional CAD designs dataset (3D-FRONT) encounters difficulties in generating scenes that align with real-world scans (Matterport) or procedurally generation (Inf3DLayout) dataset. In contrast, our method trained on the multi-source M3DLayout dataset achieves balanced performance across data types, producing more realistic and controllable layouts.
4.7 User Case Study
Figure 6: User case study results. The charts compare our method against DiffuScene and InstructScene across dining rooms, bedrooms, and living rooms. Bars represent the average user score for six metrics: Match with Text (MT), Visual Quality (VQ), Scene Richness (SR), Layout Coherence (Position) (LC-P), Layout Coherence (Orientation) (LC-O), and Overall Preference (OP).
5 Conclusion
We introduced M3DLayout, a large-scale, multi-source dataset for 3D indoor layout generation from structured text descriptions. It integrates real-world scans, professional designs, and procedurally generated scenes. Each layout is paired with structured descriptions that cover global scene summaries, relational placements of large furniture, and fine-grained arrangements of small objects. This multi-source, richly annotated structure enables the learning of diverse spatial and semantic patterns across a wide variety of indoor environments. We established a benchmark using a text-conditioned diffusion model, and experimental results validate that training on our dataset enhances the diversity and detail of generated layouts, with the Inf3DLayout subset in particular enabling more complex and richly annotated scenes. Despite these advances, our dataset has certain limitations. The structured descriptions, while comprehensive, are partly generated via language models and may contain noise. We hope M3DLayout serves as a valuable resource for advancing research in text-driven 3D scene synthesis.
Appendix A APPENDIX
A.1 Dataset Split and Implementation Details
Dataset Split. To validate the fidelity and controllability of the generated 3D layouts, we randomly split the M3DLayout dataset into 12062 layouts for training and 3018 layouts for validation. For ablation studies, models were trained on three independent datasets with corresponding splits (training/validation): 4603/1151 for 3DFront, 1347/337 for Matterport, and 6112/1530 for Inf3DLayout, respectively. Regarding the testing, the same set of text prompts ($500\times 3$), generated by GPT-4o and restricted to the bedroom, dining room, and living room, is used in all experiments to ensure fairness.
Implementation Details. The model is trained for 30k epochs using AdamW optimizer with a learning rate of $2\times 10^{-4}$ and linear noise schedule.
A.2 Baseline
We compare our method with two state-of-the-art scene generation approaches: (1) DiffuScene Tang et al. (2024), a diffusion model for 3D indoor scene synthesis that denoises unordered object attributes to produce physically plausible layouts. (2) InstructScene Lin & Mu (2024), a graph diffusion model that integrates a semantic graph with a layout decoder to synthesize 3D indoor scenes from natural language instructions. Both methods allow for the conditioning on text prompts. For inference with DiffuScene, we employ the officially released model weights for the bedroom, dining room, and living room, and follow the public implementation to train InstructScene on the same room types.
A.3 Metrics
Following prior works Tang et al. (2024); Lin & Mu (2024), we adopt Fréchet Inception Distance (FID) Heusel et al. (2017) and Kernel Inception Distance (KID) Bińkowski et al. (2018) to quantify the fidelity of scenes synthesized from layouts by measuring the similarity between generated and ground-truth top-down renderings. Meanwhile, we employ the CLIP-Score Hessel et al. (2021) to evaluate the controllability of generated layouts by computing the cosine similarity between CLIP-encoded features of the generated renderings and the given prompts. To this end, we first employ a text2mesh model Xiang et al. (2025) to generate the required object instances and then retrieve both the ground-truth and synthesized scenes conditioned on the layout. More details about self-designed object retrieval process and visualization are provided in the Appendix A.6.
A.4 Dataset Layout Visualization
We provide diverse types of 3D scene data from our M3DLayout dataset in Figure 8, which includes scenes from CAD designs sourced from the 3D-FRONT dataset at the first row, scenes derived from real-world scans, specifically from the Matterport3D dataset at the second row and procedurally generated scenes from Infinigen at the third row. These images demonstrate the flexibility of the M3DLayout dataset in representing a wide spectrum of interior environments, from synthetic CAD designs to real-world captures and generative models.
A.5 Generated Layout Visualization
We visualize more generated layouts by our model trained on the M3DLayout dataset, involving bedroom, living room, and dining room in Figure 9. From the table, it is evident that our method achieves remarkable performance in both layout coherence and the richness of objects in the generated scenes. These qualitative visualizations further highlight that our method surpasses prior state-of-the-art approaches in both fidelity and controllability.
@@ A.5 Generated Layout Visualization
A.6 Object Retrieval

### v2

@@ 3.3 Dataset Statistics and Analysis
Table 3: Quantitative comparison. Lower FID/KID ($\times$0.001) and higher Clip Score indicate better synthesis quality. FID and KID are computed with respect to the real layouts from 3D-FRONT, Matterport, and Inf3DLayout. We train InstructScene following the public implementation. The optimal result is shown in bold, and the sub-optimal result is shown with an underline.
@@
DiffuScene[tangDiffuSceneDenoisingDiffusion2024] 29.47 98.03 102.12 10.32 47.92 75.49 0.1982
InstructScene[linInstructSceneInstructionDriven3D2024] 68.58 100.54 159.27 54.70 49.23 156.62 0.1944
Ours (DIFF-M3DLayout) 57.64 87.89 70.85 36.80 34.62 50.94 0.2001
Ours (AR-M3DLayout) 87.98 107.58 57.90 57.41 53.89 41.04 0.2026
Data Composition and Source Characteristics. As detailed in Table 1, M3DLayout integrates 21,367 layouts with 433,842 object instances, averaging 20.3 objects per scene. The three data sources exhibit complementary characteristics: 3D-FRONT provides professionally designed layouts with clean structural regularity but limited small objects (0.2%); Matterport3D offers realistic scanned environments with moderate object density (12.6 objects/scene) and a balanced object distribution (39.4% small objects); while Inf3DLayout significantly enriches the dataset with high scene complexity (26.8 objects/scene) and abundant small objects (68.5%).
Comparative Advantages. As shown in Table 2, M3DLayout surpasses existing datasets in scale (21,367 scenes, 433k+ objects), layout complexity, and object size variation. Unlike datasets limited to either large furniture (L) or simple layouts, M3DLayout provides comprehensive coverage of both large and small objects (L-S) with structured textual descriptions, addressing a critical gap in current data resources for detailed scene generation.
4 Indoor Scene Generation
4.1 Problem Formulation
We formulate text-conditioned 3D indoor layout generation as a conditional generative problem. Given a scene-level natural language description $c^{\text{text}}$, the goal is to generate a structured 3D layout $x=\{o_{i}\}_{i=1}^{N}$ consisting of $N$ objects.
Each object $o_{i}$ is defined as a 3D oriented bounding box:
$o_{i}=(c_{i},x_{i},y_{i},z_{i},w_{i},h_{i},d_{i},\theta_{i}),$ (1)
where $c_{i}$ denotes the semantic class, $(x_{i},y_{i},z_{i})$ is the 3D object center, $(w_{i},h_{i},d_{i})$ are dimensions, and $\theta_{i}$ is the yaw rotation. The objective is to learn a conditional distribution $p_{\theta}(x\mid c^{\text{text}})$ that captures plausible spatial and semantic relationships among objects given the textual context.
We explore two complementary approaches for modeling this distribution: (1) a diffusion-based model, which learns to iteratively denoise noisy layout representations toward structured scenes, and (2) an autoregressive model, which generates scene objects sequentially in a structured order. Both methods are trained to reconstruct layouts consistent with the conditioning text while maintaining spatial coherence and physical plausibility.
4.2 Diffusion-based Model Architecture
Our diffusion model follows a denoising diffusion probabilistic framework similar to DiffuScene. The forward process gradually perturbs layout representations $x_{0}$ with Gaussian noise following a fixed variance schedule $\beta_{t}$, while the reverse process is parameterized by a neural network $\epsilon_{\theta}$ that predicts the noise conditioned on the text input:
The model is trained with a noise-prediction objective
$\mathcal{L}_{\text{DM}}=\mathbb{E}_{x_{0},c^{\text{text}},t,\epsilon}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t,c^{\text{text}})\|_{2}^{2}\right],$ (3)
and extra regularization terms such as an IoU-based collision penalty to encourage physically valid object arrangements. More details are in the supplementary material.
4.3 Autoregressive Model Architecture
To complement the diffusion paradigm, we further employ an autoregressive transformer model that generates scene layouts sequentially. Given the textual condition $c^{\text{text}}$, the model factorizes the layout likelihood as
$p_{\theta}(x\mid c^{\text{text}})=\prod_{i=1}^{N}p_{\theta}(o_{i}\mid o_{<i},c^{\text{text}}),$ (4)
where each object $o_{i}$ is predicted conditioned on the previously generated objects and the text.
The autoregressive model is implemented with a transformer encoder architecture. Text tokens and previously generated object embeddings form a unified input sequence for context-aware sequential prediction. The training objective minimizes the negative log-likelihood of object parameters under the predicted conditional distribution. Further details are provided in the supplementary material.
5 Experiments
5.1 Experimental Settings
We evaluate the effectiveness of autoregressive and diffusion-based methods on the proposed dataset, assessing the plausibility and controllability of generated 3D layouts. Detailed train/val splits and implementation details are in the supplementary material.
Baselines. We compare our method with two state-of-the-art scene generation approaches: DiffuScene [tangDiffuSceneDenoisingDiffusion2024] and InstructScene [linInstructSceneInstructionDriven3D2024], both of which are text-driven methods based on diffusion models. More details are provided in the supplementary material.
Metrics. Following prior works [tangDiffuSceneDenoisingDiffusion2024, linInstructSceneInstructionDriven3D2024], we adopt Fréchet Inception Distance (FID) [heusel2017gans] and Kernel Inception Distance (KID) [binkowski2018demystifying], and the CLIP-Score [hessel2021clipscore] to measure the controllability and fidelity, respectively. We employ a self-designed object retrieval method to realize instance filling from layout to scene and rendering. More details are provided in the supplementary material.
Figure 5: Density controllability in layout generation with different input texts. The first row presents input prompts for our layout generation model, showcasing variations in objects density from low to high, with minor changes in the last sentence. The second row illustrates the corresponding output results generated by our model, which adapt based on the prompt density.
5.2 Main Results
Quantitative Comparison.
The quantitative comparison results are shown in Table 3, where “DIFF-M3DLayout” and “AR-M3DLayout” represent diffusion- and autoregressive-based methods, respectively. In Table 3, the horizontal axis represents ground-truth renderings (real images) from three different datasets. The same set of 1,500 prompts is used as conditioning input for all comparative methods to generate layouts, which are then rendered as synthetic images and used together with real images to compute FID and KID. More details can be found in the supplementary material.
As demonstrated, for FID and KID, the proposed DIFF-M3DLayout drastically outperforms the state-of-the-art, achieving improvements of 10%–32% on the reference Matterport dataset and Inf3DLayout datasets, and AR-M3DLayout achieves 34%–44% improvements on Inf3DLayout dataset, which demonstrates superior generalization compared with DiffuScene and InstructScene. On 3D-FRONT, however, it falls behind DiffuScene on these metrics. This is primarily because the number of objects per scene in 3D-FRONT typically ranges from 5 to 12, whereas our method generates scenes with more than 12 objects in most cases, causing a mismatch in scene complexity distribution. As a result, the visual statistics of our generated layouts deviate more from the ground truth (3D-FRONT), which negatively impacts FID and KID. Conversely, the richer object counts and variety in the scenes generated by our model provide another advantage of our method relative to the multi-source datasets M3DLayout, which is also confirmed by the visualizations in Figure 4. Moreover, our method surpasses the baselines in CLIP-Score, demonstrating enhanced controllability and a stronger alignment between the generated scenes and the given prompts.
Qualitative Comparison. The proposed DIFF-M3DLayout is employed for qualitative comparison with diverse methods, with the results in the bedroom, dining room, and living room shown in Figure 4. Across all settings, the proposed DIFF-M3DLayout demonstrates improved semantic controllability and visual fidelity compared to Diffuscene and Instructscene. Specifically, DiffuScene can generate scenes that appear visually neat at first glance but struggles to produce small objects and exhibits limited prompt controllability, as illustrated by the living room case. InstructScene, on the other hand, fails to accurately model spatial relationships among instances, leading to disorderly object placement, as shown in the dining room case. By contrast, the proposed DIFF-M3DLayout generates precise and diverse objects in accordance with the prompt semantics—such as the small items on the shelf and six chairs in the dining room case—while effectively capturing 3D spatial arrangements, exemplified by “The bed is located near the wall corner” in the bedroom case.
We perform the ablation experiments using the diffusion-based method to validate the effectiveness of a single training dataset. More details and results are provided in the supplementary material.
5.3 Density Controllability
5.4 User Case Study
Figure 6: User case study results. The charts compare our method against DiffuScene and InstructScene across diverse rooms. Bars represent the average user score for six metrics: Match with Text (MT), Visual Quality (VQ), Scene Richness (SR), Layout Coherence (Position) (LC-P), Layout Coherence (Orientation) (LC-O), and Overall Preference (OP).
6 Conclusion
We introduced M3DLayout, a large-scale, multi-source dataset for 3D indoor layout generation from structured text descriptions. It integrates real-world scans, professional designs, and procedurally generated scenes. Each layout is paired with structured descriptions that cover global scene summaries, relational placements of large furniture, and fine-grained arrangements of small objects. This multi-source, richly annotated structure enables the learning of diverse spatial and semantic patterns across a wide variety of indoor environments. We established a benchmark using both a text-conditioned diffusion model and a text-conditioned autoregressive model, and experimental results validate that training on our dataset enhances the diversity and detail of generated layouts, with the Inf3DLayout subset in particular enabling more complex and richly annotated scenes. Despite these advances, our dataset has certain limitations. The structured descriptions, while comprehensive, are partly generated via language models and may contain noise. We hope M3DLayout serves as a valuable resource for advancing research in text-driven 3D scene synthesis.
[…]
Supplementary Material
7 Supplementary Material Overview
This supplementary document provides additional technical details and visualization results to support the main paper. The supplementary is organized into four sections:
Sec. 8 Implementation Details.
We provide detailed explanations of the dataset splitting strategy in  8.1, both diffusion and autoregressive based model details in  8.2, training details in  8.3, compared baselines details in  8.4 and metric details in  8.5.
Sec. 9 More Visualization Results.
We provide additional examples that further demonstrate the quality of our dataset, the fidelity of our generative model, and its coherence with textual inputs.
Sec. 10 Ablation.
We conduct some ablation studies to show that models trained on a single dataset overfit to its distribution and generalize poorly, whereas training on the multi-source M3DLayout dataset yields consistently balanced, realistic, and controllable layouts across diverse data types.
Sec. 11 Object Retrieval.
We present a scalable and self-developed layout-to-scene object retrieval pipeline that accurately maps generated layouts to 3D assets, facilitating reliable metric evaluation and high-fidelity visualizations.
8 Implementation Details
8.1 Dataset Split
Our experiments are conducted on a subset of the M3DLayout dataset containing 15,080 scenes. This subset is randomly divided into 12,062 layouts for training and 3,018 layouts for validation. For the ablation studies in Table 5, models are additionally trained on three independent datasets with the following splits (training/validation): 4,603/1,151 for 3D-FRONT, 1,347/337 for Matterport3D, and 6,112/1,530 for Inf3DLayout. For all evaluations, and to ensure a fair comparison with prior work, we do not use the scene description texts provided in M3DLayout. Instead, we generate 500 scene descriptions each for bedroom, dining room, and living room using GPT-4o, resulting in a total of 1,500 test descriptions. This unified test set is used across all experiments for fairness.
8.2 Model Details
Diffusion-based Model. Our diffusion-based model represents a scene as a sequence of $N$ objects, where the maximum number of objects $N$ is fixed at 120. Padding objects with a $\mathrm{PAD}$ category label is applied to maintain a consistent sequence length. The sequence is ordered by the $(x,z,y)$ coordinates of the bounding box centers, following recent best practices in the 3D part-level generation community for layout generation. Each object is parameterized by a concatenated vector $l_{i},s_{i},\cos\theta_{i},\sin\theta_{i},c_{i}],$ which includes its location ($l_{i}\in\mathbb{R}^{3}$), size ($s_{i}\in\mathbb{R}^{3}$), orientation ($\theta_{i}\in\mathbb{R}$, encoded using $\cos\theta_{i}$ and $\sin\theta_{i}$), and category label ($c_{i}\in\mathbb{R}^{C}$).
The denoiser is a UNet-based architecture built upon 1D convolutions with skip connections. It independently encodes and predicts the distinct attributes of the object. We use a pre-trained BERT encoder to extract text embeddings $z$ from the input description. This language guidance is injected into the denoiser network through cross-attention layers, enabling the network to predict the noise $\epsilon_{\phi}$ conditioned on the timestep $t$ and the text embedding $z$.
The training objective consists of the scene loss $L_{\mathrm{sce}}$, which minimizes the difference between the actual and predicted noise, supplemented with an IoU regularization term $L_{\mathrm{iou}}$ to penalize object intersections and encourage physically plausible arrangements.
Autoregressive Model. Our autoregressive model shares the same scene representation as our diffusion-based model for consistency. The autoregressive approach does not have a strict upper limit on the sequence length, but we still utilize data containing up to 120 objects per scene for training. A scene is represented as a sequence of objects, bookended by special start-of-sequence ($\mathrm{SOS}$) and end-of-sequence ($\mathrm{EOS}$) tokens.
The input text description is similarly encoded into text embeddings using a pre-trained BERT model. At each generation step, the input sequence to the model consists of the text token embeddings and the embeddings of all previously generated objects. A Transformer encoder serves as the backbone network. The attributes of each previously generated object are first projected by an MLP into an embedding space. The network then attends to this combined sequence to predict the parameters of the next object.
The training objective is the scene loss $L_{\mathrm{sce}}$, which minimizes the negative log-likelihood of the ground-truth object parameters given the input context.
8.3 Training Details.
Both our diffusion and autoregressive based models are trained for 30k epochs using the Adam optimizer. For the diffusion model, we use a learning rate of $2\times 10^{-4}$ with a step-wise decay factor of 0.5 every 10k steps and a linear noise schedule. For the autoregressive model, we use a learning rate of $1\times 10^{-4}$.
8.4 Compared Baselines
We compare our method with two state-of-the-art scene generation approaches: (1) DiffuScene [tangDiffuSceneDenoisingDiffusion2024], a diffusion model for 3D indoor scene synthesis that denoises unordered object attributes to produce physically plausible layouts. (2) InstructScene [linInstructSceneInstructionDriven3D2024], a graph diffusion model that integrates a semantic graph with a layout decoder to synthesize 3D indoor scenes from natural language instructions. Both methods allow for the conditioning on text prompts. For inference with DiffuScene, we employ the officially released model weights for the bedroom, dining room, and living room, and follow the public implementation to train InstructScene on the same room types.
8.5 Metrics
Following prior works [tangDiffuSceneDenoisingDiffusion2024, linInstructSceneInstructionDriven3D2024], we adopt Fréchet Inception Distance (FID) [heusel2017gans] and Kernel Inception Distance (KID) [binkowski2018demystifying] to quantify the fidelity of scenes synthesized from layouts by measuring the similarity between generated and ground-truth top-down renderings. Meanwhile, we employ the CLIP-Score [hessel2021clipscore] to evaluate the controllability of generated layouts by computing the cosine similarity between CLIP-encoded features of the generated renderings and the given prompts. To this end, we first employ a text2mesh model [xiang2025structured] to generate the required object instances and then retrieve both the ground-truth and synthesized scenes conditioned on the layout.
As for the calculation of FID and KID, the real images are typically taken from the training set; however, since our method and the baselines are trained on different datasets with varying data distributions, direct comparison would be unfair. To address this, we select different datasets as the source of real images, allowing for both a fair comparison and an evaluation of fidelity and controllability.
Table 4: Comparison of generation efficiency with state-of-the-art methods.
Method Generation Time per Scene (s) Generation Time per BBox (s) Average BBoxes per Scene
Diffuscene 17.4067 1.7674 9.849
InstructScene 1.3438 0.1280 10.502
Ours (DIFF-M3DLayout) 30.18 1.639 18.4
Ours (AR-M3DLayout) 0.2348 0.01587 14.8
9 More Visualization Results
9.1 Dataset Layout Visualization
We provide diverse types of 3D scene data from our M3DLayout dataset in Figure 8, which includes scenes from CAD designs sourced from the 3D-FRONT dataset at the first row, scenes derived from real-world scans, specifically from the Matterport3D dataset at the second row and procedurally generated scenes Inf3DLayout from Infinigen at the third row. These images demonstrate the flexibility of the M3DLayout dataset in representing a wide spectrum of interior environments, from synthetic CAD designs to real-world captures and generative models.
9.2 Generated Layout Visualization
We visualize more generated layouts by both our diffusion and autoregressive model trained on the M3DLayout dataset, involving bedroom, living room, and dining room in Figure 9 and Figure 10 respectively. From the table, it is evident that our method achieves remarkable performance in both layout coherence and the richness of objects in the generated scenes. We also provide more visualization to indicate strong text coherence of our generated layouts in Figure 11. These qualitative visualizations further highlight that our method surpasses prior state-of-the-art approaches in both fidelity and controllability.
10 Ablation Study
Table 5: Ablation studys of diffusion-based methods trained on different datasets. Lower FID/KID ($\times$0.001) and higher Clip Score indicate better synthesis quality. FID and KID are computed with respect to the real layouts from 3D-FRONT, Matterport, and Inf3DLayout.
Method FID ↓ KID ↓ CLIP-Score ↑
3D-FRONT Matterport Inf3DLayout 3D-FRONT Matterport Inf3DLayout
Diffusion-based Ours (3D-FRONT) 27.33 83.88 110.98 10.59 21.80 83.45 0.2083
@@ 10 Ablation Study
Autoregressive-based Ours (3D-FRONT) 30.02 85.08 117.39 12.09 24.46 95.11 0.2056
Ours (Matterport) 73.36 61.08 132.54 45.11 12.58 122.21 0.1959
Ours (Inf3DLayout) 105.24 122.19 60.73 70.81 65.21 43.21 0.2026
@@
We perform the ablation experiments using the diffusion- and autoregressive-based methods to validate the effectiveness of a single training dataset and report the results in Table 5. As shown in the table, for the diffusion-based method, when the training data and ground truth come from the same dataset, the model trained on a single dataset achieves the best FID and KID compared with the other two models. However, its performance drops significantly when evaluated on data from different datasets. This indicates that, while models trained on a single dataset can effectively fit the distribution of that dataset, they struggle to generalize to varied data. For examplprovided textual guidancee, a model trained on the professional CAD designs dataset (3D-FRONT) encounters difficulties in generating scenes that align with real-world scans (Matterport) or procedurally generation (Inf3DLayout) dataset. Furthermore, the autoregressive-based method exhibits the same characteristics. In contrast, these methods trained on the multi-source M3DLayout dataset achieves balanced performance across data types, producing more realistic and controllable layouts (see Table 3 in Section 5.2 of the present work for comprehensive data).
@@
Figure 10: Generated layouts by our autoregressive model trained on the M3DLayout dataset, visualized for randomly selected bedroom, living room, and dining room.
Figure 11: Our generated layouts exhibit strong adherence to the provided textual guidance.
11 Object Retrieval